OpenStudy (migitmack):

Really Simple Math Problem I forgot how to solve please help me. 3x+4=x

OpenStudy (anonymous):

alright do you understand the subtraction property of equality?

OpenStudy (anonymous):

dammit 1 typo

OpenStudy (migitmack):

it's not 1

OpenStudy (anonymous):

3x + 4 = x 4 = -3x + x 4 = -2x x = -2

OpenStudy (anonymous):

3x + 4 = x 3x-3x+4 = -3x + x [Subtraction property of equality says that if you subtract one thing on one side of the equation you must subtract it on the other side. In this case Josh is trying to move the x's to one side so he subtracts 3x from both sides] 4 = -2x [He adds like terms. -3x + x = (-3+1)x=-2x] x = -1 [Lastly he used the division property of equality which says that if you divide one side by a number, you must divide the other side by the same number. He does this because he wants to get x by itself so he divides both sides by -2]

OpenStudy (anonymous):

lol i solved this question last year but cant rememebr the answer i think is 2.

OpenStudy (anonymous):

really nicely explained, Outkast3r! :)